OpenStudy (anonymous):

Help with Algebra 2??

OpenStudy (solomonzelman):

yes, what do you need help?

OpenStudy (solomonzelman):

\(\large\color{slate}{ x=\sqrt{3x+10} }\) to start, raise each side to the second power.

OpenStudy (solomonzelman):

\(\large\color{slate}{ \color{red}{(}x\color{red}{)^2}=\color{red}{(}\sqrt{3x+10}\color{red}{)^2} }\)

OpenStudy (solomonzelman):

tell me, what do you get after doing this?

OpenStudy (anonymous):

x^2=3x+10?

OpenStudy (solomonzelman):

yes.

OpenStudy (solomonzelman):

Now, I would subtract 3x, and then 10, from both sides. Then I would factor both sides.

OpenStudy (solomonzelman):

I mean factor the left side, not both sides. Sorry.

OpenStudy (solomonzelman):

\(\large\color{slate}{ x^2=3x+10 }\) will do a shortcut, subtracting the entire \(\large\color{slate}{ \color{blue}{3x+10} }\) from both sides, \(\large\color{slate}{ x^2\color{blue}{-3x-10}=3x+10 \color{blue}{-3x-10}}\)

OpenStudy (solomonzelman):

what do you get after doing this?

OpenStudy (solomonzelman):

lost?`

OpenStudy (solomonzelman):

nello, if you are lost please reply, don't hesitate....

OpenStudy (anonymous):

yeah... sorry

OpenStudy (solomonzelman):

Sure, sure.... no need for being sorry. Lets see, do you understand how we got till \(\large\color{slate}{ x^2=3x+10 }\) ?

OpenStudy (anonymous):

we squared each side?

OpenStudy (solomonzelman):

yes, exactly!

OpenStudy (solomonzelman):

Now, lets go ahead and subtract 3x from both sides. \(\large\color{slate}{ x^2\color{blue}{-3x}=3x+10 \color{blue}{-3x}}\) can you tell me what we get after doing this?

OpenStudy (anonymous):

now what do I do? should I get the x on one side?

OpenStudy (solomonzelman):

tell me what cancels on the right side (if anything) ?

OpenStudy (anonymous):

the 10

OpenStudy (solomonzelman):

the 10 stands there

OpenStudy (solomonzelman):

\(\large\color{slate}{ x^2\color{blue}{-3x}=\color{black}{\cancel {3x}}+10 \color{blue}{\cancel{-3x}}}\)

OpenStudy (anonymous):

then the 3?

OpenStudy (solomonzelman):

3x, you mean?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

so now its x^-3x=10

OpenStudy (anonymous):

x^2

OpenStudy (solomonzelman):

yes, \(\large\color{slate}{ x^2-3x=10 }\)

OpenStudy (solomonzelman):

\(\large\color{slate}{ x^2-3x=10 }\) subtract 10 from both sides, you get? \(\large\color{slate}{ x^2-3x\color{blue}{-10}=10 \color{blue}{-10}}\)

OpenStudy (anonymous):

If I can get 10 to the left too I can factor them and the answer will be 5 and -2?

OpenStudy (solomonzelman):

yes, very good x=-2 or x=5.

OpenStudy (solomonzelman):

but we must check them

OpenStudy (solomonzelman):

\(\large\color{slate}{ x=\sqrt{3x+10}}\) True/False: 1. \(\large\color{slate}{ 5=\sqrt{3(5)+10}}\) 2. \(\large\color{slate}{ -2=\sqrt{3(-2)+10}}\)

OpenStudy (solomonzelman):

can you answer this question?